### 12 Dec 2023 v1

# Variety and Quality over Quantity: Towards Versatile Instruction Curation

Abstract
Instruction fine-tuning, involving the refinement of pre-trained LLMs using datasets accompanied by natural instructions, is a powerful approach. However, its effectiveness is hindered by the redundancy and deficiencies in LLM-generated instruction datasets. In this paper, we introduce a highly effective and versatile paradigm for selecting diverse and high-quality instruction-following data from fine-tuning datasets. We first employ the dataset enhancement and expansion to augment the dataset with more diverse and high-quality data, then we apply variety compression and quality compression sequentially to curate the desired dataset. Our experimental results showcase that, even with a limited quantity of high-quality instruction data, LLMs consistently maintain robust performance across both natural language understanding tasks and code generation tasks. Notably, they outperform models trained on significantly larger instruction datasets in certain instances.
1 Introduction
In recent years, Large Language Models (LLMs) have gained prominence for their remarkable effectiveness in natural language comprehension tasks (OpenAI, 2023; Yang et al., 2023). High-quality pretrained LLMs like LLaMA2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023) and StarCoder (Li et al., 2023b) are readily available, facilitating their customization for versatile applications. One popular fine-tuning approach, known as instruction tuning (Wei et al., 2022; Ouyang et al., 2022), involves fine-tuning pre-trained LLMs using datasets accompanied by natural language instructions. Its relative simplicity and affordability make it a preferred method for improving language models.
Figure 1: Proposed Instruction Dataset Curation Paradigm Overview
However, a significant challenge in the instruction tuning process arises from the observation that fine-tuning with larger instruction datasets does not always guarantee better results, yet demanding more computational resources. This issue stems from the redundancy and deficiencies in LLM-generated instruction datasets (Luo et al., 2023). The LLM-generated dataset, proposed as a substitute for resource-intensive manual data selection, lacks supervision regarding the diversity and quality of the generated data. Dataset redundancy poses the risk of the model overfitting to the fine-tuning data, thereby hindering its capacity for generalization. Concurrently, the inclusion of low-quality or potentially harmful data can degrade the performance of LLMs.
The concern surrounding the instruction fine-tuning process has prompted researchers to explore methods for fine-tuning large models using a limited quantity of high-quality data in order to attain robust performance. Encouragingly, several recent studies have validated the efficacy of this approach, demonstrate that large language models can achieve impressive results even when provided with a limited quantity of high-quality instruction-following data (Zhou et al., 2023; Wei et al., 2023). However, current research is still primarily focused on limited models and tasks. There is still a lack of systematic exploration on compression principles and techniques.
Inspired by the previous work on curating fine-tuning dataset for LLMs, we propose a new LLM instruction dataset curation paradigm, presented in figure 1. This paradigm is designed to generate a diverse and high-quality dataset for fine-tuning LLMs, while significantly reduce the quantity of the dataset. Our paradigm is composed of three steps. First, we employ the "Dataset Enhancement and Expansion", enhancing the accuracy of the original instructions’ answers and introduce new high-quality data. Subsequently, the "Variety Compress" approach eliminates redundancy from both the original and newly generated data. Finally, the "Quality Compress" process assesses the quality of each item in the compressed data, selecting high-quality ones to form the final curated dataset.
To validate the effectiveness of our paradigm, we employ the final curated instruction datasets for fine-tuning open-source LLMs. We conduct extensive experiments to evaluate these fine-tuned LLMs’ performance in both natural language understanding (NLU) tasks and code generation tasks, considering 6 metrics. Results consistently show that LLMs maintain robust performance even with a limited quantity of high-quality instruction data. In certain cases, they even surpass models trained on larger datasets.
Our contribution can be summarized as follows:
1. We propose a highly effective and versatile paradigm for the selection of diverse and high-quality instruction-following data from fine-tuning datasets.
2. Our experiment results demonstrate that with a small quantity of instruction data selected by our paradigm, LLMs can still maintain robust performance, and in some cases, even outperform models trained with larger instruction datasets.
3. Our findings provide valuable insights, suggesting that instead of indiscriminately feeding large volumes of data into LLMs during training, selecting a smaller quantity of high-quality, multifaceted data for fine-tuning not only conserves time and resources but also preserves and, in some tasks, enhances performance.
2 Related Works
2.1 Instruction Tuning
While LLMs have demonstrated cross-task generalization capabilities, they often fall short in aligning with users’ prompts when applied to real-world user communications. To overcome this limitation, OpenAI enlisted professional human annotators to generate an extensive corpus with human instructions. Leveraging this instruction dataset, OpenAI fine-tuned its GPT-3 model to create InstructGPT (Ouyang et al., 2022). Subsequently, recent research has pivoted towards automatic instruction generation frameworks. Alpaca (Taori et al., 2023) proposed adopting the self-instruct method by utilizing ChatGPT to generate data for fine-tuning. Vicuna (Chiang et al., 2023) utilized data collected from ShareGPT.com, a platform where users share their conversations with ChatGPT, for fine-tuning their models. WizardLM (Xu et al., 2023) and WizardCoder (Luo et al., 2023) introduced the Evol-Instruct method, which involves evolving existing instruction data to generate more diverse and complex data.
LLM-generated datasets often fall short in dataset variety and quality, due to limited measures to guarantee and control the output of LLMs. This hidden issue further constrains the potential for performance improvement in LLMs.
2.2 Instruction Dataset Curation
To enhance the model’s performance following instruction fine-tuning, some researchers have concentrated on filtering out low-quality data during the fine-tuning stage. LIMA (Zhou et al., 2023) demonstrates that fine-tuning a robust pre-trained language model on 1000 high-quality, human-curated examples can yield remarkable and competitive results. Instruction Mining (Cao et al., 2023) introduces a linear rule for selecting high-quality instruction data, eliminating the need for human annotation, as seen in LIMA. Du et al. present a model-oriented data selection (MoDS) approach, which selects instruction data based on a new criteria considering three aspects: quality, coverage and necessity. Li et al. introduce a self-guided methodology for LLMs to autonomously discern and select cherry samples from vast open-source datasets, effectively minimizing manual curation and potential cost.
In contrast to the aforementioned works, which focused solely on a single LLM task, our paradigm offers a versatile solution for effectively selecting diverse and high-quality data from the raw dataset used for fine-tuning, encompassing both NLU tasks and code-related tasks.
3 Methodology
As shown in figure 1, our paradigm consists of three steps: Data Enhancement and Expansion, Variety Compress and Quality Compress. In this section, we’ll introduce these three steps in details.
3.1 Data Enhancement and Expansion
The original dataset often contains brief and straightforward prompts, resulting in simplistic responses lacking significant semantic information. These instances of low-quality data are part of the original dataset, suggesting that relying solely on compressing it would yield limited performance improvements. To boost the quality of our compressed dataset, we find data enhancement and expansion necessary. The workflow of these two phase is presented in figure 2.
Figure 2: Instruction Dataset Enhancement and Expansion
3.1.1 Data Enhancement
In data enhancement phase, we replace the original dataset’s output section with GPT-4’s response. The output section of the original instruction dataset is often generated by humans or other LLMs. While generally accurate and aligned with users’ intentions, these outputs frequently lack comprehensive explanations due to brevity. For instance, a prompt like "Write a function to generate the nth Fibonacci number" might yield a concise Python code answer in the original dataset, lacking detailed coding logic and considerations for boundary cases.
Recognizing GPT-4’s exceptional natural language and code comprehension, we choose to utilize GPT-4 to re-generate the "output" section in the dataset. The prompt given to GPT-4 comprises the original dataset’s "instruction" and "input" sections. Compared to the original output, GPT-4 tends to generate more detailed answers, including comments for programming requests and additional reasoning steps to address logic questions.
3.1.2 Data Expansion
Inspired by the instruction rewrite method proposed in WizardLM and WizardCoder (Xu et al., 2023; Luo et al., 2023), our approach is similarly dedicated to producing instructions that are diverse and intricate. We implement the "Instruction-Elevate" process, where we instruct GPT-4 to function as a prompt re-writer, generating challenging instructions based on specified generation rules.
Given the variation in instruction dataset content for Natural Language Understanding (NLU) and code generation tasks, we provide distinct settings for GPT prompts to increase complexity. Additional details are available in Appendix A.
After generating new instructions and inputs, we use GPT-4 to produce corresponding outputs, concluding an expansion round. Each new instruction item corresponds to an original one, resulting in a dataset of the same size as the original after each round. We iterate this process for $k$ rounds, combining the expanded datasets with the enhanced dataset to create a final expanded dataset.
3.2 Variety Compress
The "Variety Compress" step aims to eliminate duplicated data from both the original dataset and the expanded dataset, generated in the previous step. Duplicated items may lead LLMs to overfit on specific topics, reducing their ability to generalize knowledge. The workflow is depicted in Figure 3. To identify duplication, we opt to analyze embeddings of the expanded dataset.
Initially, GPT generates embeddings with 1536 dimensions for each item, but this proves unwieldy for analysis. To address this, we aim to reduce the embedding dimension and devise a method to represent data differentiation.
Figure 3: Variety Compress
3.2.1 Normalization & Dimension Reduction
We first calculate the covariance matrix of the given features, perform eigenvalue decomposition on the covariance matrix to obtain eigenvalues and eigenvectors, then choose top $k$ eigenvectors corresponding to the largest $k$ eigenvalues, where $k$ is the target reduced dimension. It’s essential to acknowledge that as we compress the dimension of feature vectors, some information loss is inevitable. Careful consideration is needed when selecting the target dimension denoted as $k$ to ensure retention of the majority of the original information.
Before directly reducing the embedding data’s dimension, feature normalization is necessary. Given covariance matrix’s sensitivity to varying feature scales, normalization becomes crucial as larger-scale features may dominate the covariance matrix, impacting the computation of principal components and potentially introducing bias. Normalization mitigates this bias, ensuring equal contribution from each feature. After normalization, we undertake the following steps to ascertain the target reduced dimension $k$. We designate $k$ as the reduced dimension, aiming for a value where $k$ features can retain approximately 95% of the variance present in the original embeddings, thus preserving a substantial amount of information.
3.2.2 Row Variance for Dimension-reduced Features
Following dimension reduction, data undergoes linear transformation and projects into a new coordinate system defined by selected $k$ features. Reduced-dimensional data corresponds to this projection, allowing analysis of row variance to identify items with significant differences. Row variance measures variability in the reduced space across all $k$ dimensions for each data point. High variance suggests substantial positional changes along that dimension, aiding in identifying diverse data points. After calculating row variances, we select items with the highest 20% row variances to construct the variety-compressed dataset, ensuring diverse instructions are chosen.
3.3 Quality Compress
Figure 4: Quality Compress
"Quality Compress" is designed to identify high-quality instruction data from a compressed dataset. High-quality instructions and their outputs should meet specific criteria: clear and detailed instructions, presenting a reasonable level of complexity for advanced LLMs. The expected output should align accurately with instructions and provide comprehensive explanations.
Rating instruction data quality is challenging without official quantitative metrics. Due to the dataset’s size and associated costs, enlisting professional annotators for individual scoring is unrealistic. Therefore, we leverage GPT-4, an advanced LLM, as an instruction scorer, generating GPT quality score for each instruction. Furthermore, recognizing that the total length of instruction data indicates information richness, we incorporate this into the quality score calculation. Unlike other metrics, length is quantitatively analyzable. A positively correlated mapping function derives a lengthwise semantic score based on the instruction data’s length. Combining the GPT quality score and lengthwise semantic score produces the final quality score. Instruction items with high-quality scores compose the final quality-compressed dataset, as illustrated in Figure 4.
3.3.1 GPT Quality Evaluation
We engage GPT-4 to evaluate the instruction data across four dimensions: accuracy, explanation, clarity, and difficulty, each contributing to a distinct portion of the total GPT score.
Accuracy (35%): Reflects the correctness of the output solution concerning the instruction’s problem. A high score is awarded for a solution that perfectly addresses all requirements.
Explanation (25%): Applied to the output, this metric rewards detailed explanations, such as comments in coding tasks and reasoning steps within a chain of thoughts.
Clarity (15%). Assesses how well the instruction articulates requests, emphasizing clear details and avoiding ambiguity or overly general questions.
Difficulty (25%). Focuses on the complexity of the instruction’s problem. Higher scores are given for requests with reasoning constraints and additional depth of thought.
These metrics’ proportions are assigned based on their contributions to overall instruction quality. The template guiding the GPT-4 scores is presented in Appendix B.
In our practical experience, we observed that GPT-4 consistently assigns high total scores (over 90) to all instruction data, even when explicitly instructed not to do so. This behavior stems from GPT-4’s training as an overall friendly and positive agent, posing a challenge by reducing differentiation in GPT quality scores. To address this, we implement the following steps for more reasonable and differentiated scores:
Require scores along with detailed explanations. Instructing GPT-4 to provide a comprehensive rationale along with a score yields more reasonable results, preventing indiscriminate high-level assessments. While we may not use these explicit explanations, mandating GPT-4 to articulate its reasoning offers an additional self-checking opportunity.
Provide few-shots scoring examples. Before scoring instruction items, we present manually scored examples as guidelines. Offering three examples with scores representing poor, average, and high quality helps GPT-4 recognize low-quality data and understand how to appropriately score it.
3.3.2 Lengthwise Semantic Evaluation
The lengthwise semantic complexity score is determined by a linear mapping function. Instructions with the shortest total length, including both the instruction and output, receive the lowest score, while those surpassing a specified threshold attain the highest score.
Combining the GPT quality assessment scores and lengthwise semantic scores produces the final quality score, facilitating the selection of high-quality items. In Appendix C, we present an overview of the total quality score distributions across both tasks. Our meticulous efforts ensure substantial differentiation in quality scores, allowing precise identification and selection of high-quality data.
4 Experiments
To validate the effectiveness of our paradigm, we apply our method to two extensively studied tasks: Code Generation tasks and NLU tasks. Subsequently, we conduct comprehensive experiments to evaluate the performance of our approach.
4.1 Experiments Setup
4.1.1 Basic Foundation Models and Base Datasets
For code generation tasks, we leverage StarCoder 15B (Li et al., 2023b), a widely-used large code language model trained on diverse sources, including 80+ programming languages, Git commits, GitHub issues, and Jupyter notebooks. Our base dataset, Code Alpaca (Chaudhary, 2023), consists of 20k instruction-following code instances for fine-tuning Code LLMs.
In NLU tasks, we employ the state-of-the-art model Mistral 7B (Jiang et al., 2023), known for superior performance compared to other 7B models and outperforming larger models in certain benchmarks. Our base dataset for NLU tasks is the Open Platypus dataset (Lee et al., 2023), comprising 25k curated examples focused on enhancing LLMs’ STEM and logic knowledge. This dataset, composed of 11 open-source datasets, mainly contains human-designed questions, with only 10% generated by an LLM.
The instruction tuning implementation details for both tasks can be found in Appendix D
4.2 Benchmarks and Metrics
We have selected six widely-used benchmarks across the two tasks. For code generation tasks, we incorporate HumanEval and MBPP. And for NLU tasks, our selected benchmarks include HellaSwag, ARC Challenge, TruthfulQA, and MMLU. The details of these selected benchmarks can be found in Appendix E.1.
In code generation tasks, our metric of choice is pass@k, defined in the same manner as by Chen et al.. The formula for calculating pass@k is presented as:
$pass@k:=\mathbb{E}_{problems}[1-\frac{C(n-c,k)}{C(n,k)}]$
Here, $n$ represents the number of generated answers for each question, and $c$ denotes the number of correct answers for each question. In our experiments, we choose pass@1 as the metric.
For NLU tasks, we opt for accuracy as the metric, aligning with the approach adopted by other researchers.
4.3 Experiment Results
We compare models fine-tuned on our paradigm’s final curated dataset with other state-of-the-art (SOTA) pretrained LLMs and instruction-tuned LLMs across both tasks. The details of the selected models for comparison is presented in Appendix E.2
4.3.1 Code Generation Tasks
As depicted in Table 1, our paradigm’s fine-tuned model outperforms most of the models in code generation tasks. While our model is nearly 2% lower than the current state-of-the-art 15B model, WizardCoder, on both benchmarks, it’s important to note that our paradigm utilizes only about one-eighth of the instruction data used by WizardCoder. Considering the size of the instruction dataset, our paradigm demonstrates robust performance, showcasing the ability to achieve performance close to the state-of-the-art with a significantly smaller amount of data.
Table 1: LLMs Performance Comparison in Code Generation Tasks
Model Fine-tuning Data Size HumanEval MBPP
CodeT5+ 16B Pretrained 0.309 -
Code LLaMA-13B 0.360 0.470
StarCoder-15B 0.336 0.436
[…]
Code LLaMA Instruct-13B approx. 5B tokens 0.427 0.494
WizardCoder-15B 78k instructions 0.573 0.518
[…]
Ours-15B 10k instructions 0.550 0.495
4.3.2 Natural Language Understanding Tasks
Table 2: LLMs Performance Comparison in NLU Tasks
Model Fine-tuning Data Size HellaSwag ARC TruthfulQA MMLU
LLaMA-7B Pretrained 0.778 0.509 0.343 0.357
LLaMA-13B 0.809 0.561 0.395 0.476
LLaMA2-7B 0.771 0.432 0.333 0.444
LLaMA2-13B 0.807 0.488 0.419 0.556
Mistral-7B 0.823 0.602 0.426 0.627
[…]
Vicuna-13B 70k conversations 0.801 0.530 0.518 0.513
WizardLM-7B 70k instructions 0.771 0.516 0.447 0.427
WizardLM-13B 70k instructions 0.777 0.572 0.505 0.523
Platypus2-13B 25k instructions 0.826 0.613 0.449 0.567
Camel-Platypus2-13B 25k instructions 0.836 0.608 0.496 0.565
Stable-Platypus2-13B 25k instructions 0.822 0.627 0.525 0.583
[…]
Ours-7B 15k instructions 0.844 0.643 0.490 0.645
Table 2 presents the NLU task comparison results. Notably, our final instruction-tuned model consistently outperforms other 7B models on all benchmarks. Even among 13B models, our model outperforms in all benchmarks except TruthfulQA. With only 7 billion parameters and 15k instructions, significantly fewer than other instruction-tuned models, our model achieves the highest average benchmark score at 0.656.
We also compared our paradigm’s final curated dataset with a randomly selected dataset of the same size in both tasks. The results demonstrate that merely reducing the dataset quantity, without accounting for the diversity and quality of the data, does not lead to performance improvement. These experiments affirm our paradigm’s versatile effectiveness in code generation and NLU tasks. The paradigm excels in generating diverse, high-quality data, leveraging it in the instruction-tuning process to achieve SOTA or near-SOTA performance.
4.4 Paradigm Ablation Experiments Results
Our paradigm ablation experiment begins with the original base dataset serving as the input for our paradigm. Subsequently, we generate the expanded dataset, variety-compressed dataset, and the quality-compressed dataset. These datasets are then utilized for fine-tuning the basic foundation models. We assess the benchmark performance of these models to validate the effectiveness of each component of our paradigm.
4.4.1 Code Generation Tasks
Table 3 provides an overview of the paradigm experiments conducted on code generation tasks. For data expansion, we repeatedly perform the expansion step two times, resulting in a 60k size instruction dataset.
Table 3: Paradigm Ablation Experiment Results (Pass@1) in Code Generation Tasks
Dataset Type Size HumanEval MBPP
Base Dataset 20k 0.4091 0.4662
Data Expansion 60k 0.5342 0.4874
Variety Compress 12k 0.5412 0.4887
Quality Compress 10k 0.5503 0.4949
The table illustrates our paradigm leads to a significant enhancement in the performance of the fine-tuned model across both benchmarks. Notably, our final compressed dataset, although roughly half the size of the original dataset, outperforms the latter by nearly 15% on the HumanEval and 3% on the MBPP. Furthermore, we observe a consistent improvement in model performance on both benchmarks after each step of the paradigm. This implies that the instruction’s quality is steadily increasing at each stage. These results strongly affirm that, compared to the quantity of the instruction dataset, the variety and quality of the dataset play a more significant role in enhancing performance.
4.4.2 Natural Language Understanding Tasks
Table 4: Paradigm Ablation Experiment Results in NLU Tasks
Dataset Type Dataset Size HellaSwag ARC Challenge TruthfulQA MMLU
Base Platypus Dataset 25k 0.82788 0.61543 0.44481 0.62619
Data Expansion 100k 0.83308 0.62372 0.44718 0.63065
Variety Compress 20k 0.83947 0.63311 0.45615 0.64199
Quality Compress 15k 0.84415 0.64334 0.48985 0.64519
Table 4 presents our paradigm experiment results on four NLU benchmarks. For data expansion, we iteratively perform the expansion step three times, resulting in a 100k size instruction dataset.
The table results affirm our paradigm’s effectiveness in NLU. Despite a reduction in size by 10k instances compared to the original dataset, our final compressed dataset maintains robust performance, showing improvements ranging from nearly 2% to 4% on each benchmark. The observed performance improvement pattern in code generation tasks also applies to NLU tasks, with each step enhancing data quality.
It’s crucial to note that unlike the Code Alpaca (Chaudhary, 2023) dataset, the Open Platypus (Lee et al., 2023) dataset for NLU tasks is already carefully curated. The results for this dataset underscore that our paradigm is effective not only for LLM-generated datasets but also in elevating the quality of already high-quality datasets, contributing to improved fine-tuned model performance while reducing the dataset size.
5 Discussions
5.1 Composition of The Final Curated Dataset
(a) Source of the final curated 10k dataset in code generation tasks
(b) Source of the final curated 15k dataset in NLU tasks
Figure 5: Composition of The Final Curated Dataset
We take a step further to analyze the composition of the final curated dataset, unraveling the origins of diverse and high-quality instruction items. Figure 5 presents the source proportions of the final curated dataset for code generation and NLU tasks, yielding several noteworthy conclusions.
For LLM-generated instruction datasets like Code Alpaca, only a small proportion of the final dataset emanates from the original dataset (Fig. 4(a)). The majority of high-quality data is derived from our paradigm’s first step—the expanded dataset. This emphasizes our paradigm’s significant role in generating a diverse and high-quality dataset, especially for datasets without meticulous curation.
In contrast, for a curated and high-quality instruction dataset like Open Platypus, the portion of the original dataset in the final dataset increases (Fig. 4(b)). The proportions of the final curated dataset in NLU tasks reveal an almost equal distribution among the four sub-datasets, demonstrating that even for an initially high-quality dataset, our paradigm excels in generating and selecting numerous high-quality data points based on the original dataset.
These conclusions affirm that the bulk of the final dataset primarily comprises data from the expanded dataset. While proportions of original dataset data contributing to the final dataset may vary based on the original dataset’s quality, our paradigm consistently showcases its ability to extract high-quality segments from the original dataset and augment them with diverse and high-quality instructions.
5.2 GPU Hours and Carbon Emission
By compressing the size of the instruction dataset, we aim to reduce the GPU hours required for instruction tuning, contributing to a subsequent decrease in carbon emissions. Table 5 illustrates the impact of different dataset sizes on GPU hours and CO2 emissions. GPU hours are calculated under the same settings of training epoch and batch size, while carbon emissions are computed using an online machine learning CO2 calculator [^1] https://mlco2.github.io/impact/#co2eq.
Table 5: Analysis of GPU hours and Carbon Emission with different dataset size. GPU hours in the table are measured in hour, CO2 emission is in kg CO2 eq.
Dataset Size GPU Hours CO2 Emission
Code Generation Tasks
Original 20k 50.82 4.58
Expanded 60k 185.6 16.7
Curated 10k 31.6 2.84
NLU Tasks
Original 25k 40.24 3.62
Expanded 100k 149.76 13.48
Curated 15k 23.71 2.13
The table shows a substantial reduction in GPU hours and lower carbon emissions when fine-tuning with the final curated dataset. This comparison demonstrates that our paradigm not only accelerates fine-tuning but also promotes environmental sustainability while maintaining robust high performance.
5.3 Limitations
Redundancy in Data Expansion: The first step involves data expansion, introducing some data redundancy as the new instructions are primarily based on rewriting the original ones. Despite mitigation in the Variety Compress step, the impact of duplicated instructions from the expansion phase should be considered.
Subjectivity in Quality Evaluation: Quality evaluation relies heavily on the GPT Quality Evaluation, and despite careful criteria, additional statistical analysis beyond the length factor could enhance the precision of high-quality data selection.
Unexplored Compression Limits: Our study does not explicitly investigate the compression limits of the original dataset. Understanding compression limits for both datasets is essential for optimizing the trade-off between size and performance.
5.4 Future Works
Diverse Instruction Evolution Strategies: We aim to explore strategies for expanding data that are less reliant on the original dataset, utilizing techniques such as prompt engineering and implementing stringent data supervision strategies for generating diverse instructions.
Comprehensive Metrics for Quality Evaluation: Future work will involve enhancing quality evaluation metrics beyond the current focus on instruction length. This includes integrating more comprehensive metrics like information entropy analysis, coherence analysis, and word frequency analysis to provide a nuanced assessment of instruction quality.
In-Depth Exploration of Compression Limits: Further investigation into the compression limits of the original instruction dataset is also a focus. We will delve into understanding whether the observed robust performance extends to smaller curated datasets and explore the detailed relationship between compression ratios and instruction datasets.
6 Conclusions
This paper introduces a novel paradigm for curating large language model instruction datasets to enhance the quality and diversity of data for fine-tuning LLMs. The paradigm unfolds in three steps: Data Enhancement and Expansion, Variety Compress, and Quality Compress.
Experimental results shows that our fine-tuned models consistently achieve either SOTA or nearly SOTA performance across both code generation and natural language understanding tasks. These experiments affirm the paradigm’s versatile effectiveness, demonstrating its ability to generate diverse and high-quality data. Integrating this curated data into the instruction-tuning process empowers LLMs to achieve superior performance across various benchmarks.
Appendix A GPT Prompt Templates For Data Expansion
We adopt two different prompt settings for natural language understanding tasks and code generation tasks.
A.1 Natural Language Understanding Tasks
{spverbatim}
SYSTEM MESSAGE: I want you act as a professional prompt re-writer. USER PROMPTS: Your objective is to rewrite a given prompt into a more complex version using data format to make those famous AI systems more difficult to handle. But the rewritten prompt must be reasonable and must be understood and responded by humans. You can increase the difficulty using, but not limited to, the following methods: (1) The depth and breadth of the inquiry can be increased. (2) Replace general concepts with more specific concepts. (3) If original problem can be solved with just a few simple thinking processes, you can rewrite it to explicitly request multiple-step reasoning.
#Instruction# Instruction #Input# Input
A.2 Code Generation Tasks
{spverbatim}
SYSTEM MESSAGE: I want you act as a professional Prompt Rewriter. USER PROMPTS: Please increase the difficulty of the given programming test question a bit. You can increase the difficulty using, but not limited to, the following methods: (1) Add new constraints and requirements to the original problem, adding approximately 10 additional words. (2) Replace a commonly used requirement in the programming task with a less common and more specific one. (3) If the original problem can be solved with only a few logical steps, please add more reasoning steps. (4) Provide a piece of erroneous code as a reference to increase misdirection. (5) Propose higher time or space complexity requirements, but please refrain from doing so frequently.
#Instruction# Instruction #Input# Input
Appendix B GPT-4 Score Template
{spverbatim}
SYSTEM MESSAGE: We would like to request your feedback on the performance of an AI assistant. The assistant provides outputs for instruction and input (if any). USER PROMPTS: Please score the response to the instruction and input according to the following criteria. The maximum score is 100 points, and it consists of 4 parts: 1. Clarity (15 points): Assign a score based on how effectively the instruction conveys the problem. High-quality, clear questions score higher. 2. Difficulty (25 points): Rate the complexity of the instruction’s problem. Higher difficulty should receive a higher score. 3. Explanations (25 points): Assess if the response includes detailed explanations alongside any code provided. The more comprehensive the explanation, the higher the score. 4. Accuracy (35 points): Score the response based on the accuracy and correctness of the solution to the instruction’s problem. Higher accuracy should receive a higher score. Here’s some examples and socres you can follow: ### Example 1: ### Instruction: EXAMPLE INSTRUCTION 1 ### Response: EXAMPLE OUTPUT 1 ### Score for Example 1: SCORE 1 ### Example 2: ### Instruction: EXAMPLE INSTRUCTION 2 ### Input: EXAMPLE INPUT 2 ### Response: EXAMPLE OUTPUT 2 ### Score for Example 2: SCORE 2 ### Example 3: ### Instruction: EXAMPLE INSTRUCTION 3 ### Response: EXAMPLE OUTPUT 3 ### Score for Example 3: SCORE 3
Please score the upcoming Instruction, Input and Response based on these examples across four dimensions, and then add the four scores together to get the total score. Try to avoid getting a full score as much as possible. Please first output a single line containing the total score number only. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias. ### Instruction: INSTRUCTION ### Input: INPUT ### Response: OUTPUT
Appendix C Quality Score Distribution
We have gathered the quality scores for the variety compressed dataset following the second step of our paradigm, both in code generation tasks and NLU tasks. The score distributions are depicted in Figure 6. Notably, the quality scores exhibit an approximately normal distribution within the score interval of 60 to 100 for both tasks. This observation validates the effectiveness of our scoring strategies in discerning low-quality data. It should be noted that the minor bumps near 0 stem from connection errors or OpenAI API calling ratio constraints, resulting in GPT scores of 0 for certain instructions.
(a) Quality Score Distribution in Code Generation Tasks (12k Data)
(b) Quality Score Distribution in NLU Tasks (20k Data)
Figure 6: Quality Score Distribution
Appendix D Experiments Implementation Details
For both foundation models, we conduct training on Azure Machine Learning Studio’s cluster, utilizing 4 nodes, each equipped with 8 V100 GPUs featuring DeepSpeed Zero-3 (Rajbhandari et al., 2019) offload. Specifically, during the fine-tuning of Mistral 7B, we employ LoRA (Hu et al., 2021). This strategy is chosen for its ability to ensure a more steady convergence of loss, resulting in better performance. The detailed fine-tuning arguments are outlined in Table 6.
Table 6: Fine-tuning Arguments for StarCoder 15B and Mistral 7B
Arguments StarCoder Mistral
model_max_length 1024 2048
batch_size 8 8
num_epoch 3 3
learning_rate 2e-5 2e-5
fp16 True True
lora_r - 16
lora_alpha - 32
lora_dropout - 0.05
Appendix E Benchmarks and Compared LLMs
E.1 Benchmarks
Below is an introduction to these chosen benchmarks:
  • HumanEval (Chen et al., 2021). HumanEval is utilized to gauge functional correctness in synthesizing programs from docstrings. Comprising 164 original programming problems, it assesses language comprehension, algorithms, and simple mathematics.
  • MBPP (Austin et al., 2021). The MBPP (Mostly Basic Python Problems) dataset consists of around 1,000 crowd-sourced Python programming problems. These are designed to be solvable by entry-level programmers, covering programming fundamentals and standard library functionality. In our experiments, to align with others, we select 400 questions.
  • HellaSwag (Zellers et al., 2019). HellaSwag is a challenge dataset containing 70k multiple-choice questions for evaluating commonsense Natural Language Inference (NLI). While its questions may be trivial for humans (>95% accuracy), they pose a challenge for state-of-the-art models.
  • ARC Challenge (Clark et al., 2018). The AI2’s Reasoning Challenge (ARC) dataset is a multiple-choice question-answering dataset containing questions from science exams ranging from grade 3 to grade 9. It is split into two partitions: Easy and Challenge. The Challenge partition consists of 25k questions that require reasoning.
  • TruthfulQA (Lin et al., 2022). TruthfulQA is a benchmark designed to measure whether a language model is truthful in generating answers to questions. The benchmark comprises 817 questions spanning 38 categories. Questions are crafted so that some humans might answer falsely due to false beliefs or misconceptions.
  • MMLU (Hendrycks et al., 2020). MMLU (Massive Multitask Language Understanding) is a new benchmark intended to measure knowledge acquired during pretraining. It evaluates models exclusively in zero-shot and few-shot settings, making it more challenging and akin to human evaluation. The benchmark covers 57 subjects across STEM, humanities, social sciences, and more, ranging in difficulty from elementary to advanced professional levels, testing both world knowledge and problem-solving ability.
E.2 Compared LLMs
The selected models for comparison in code generation tasks include:
  • CodeT5+ & InstructionCodeT5+ (Wang et al., 2023). CodeT5+ is a new family of open code LLMs with an encoder-decoder architecture trained on various pretraining tasks. InstructionCodeT5+ is further fine-tuned on the Code Alpaca dataset.
  • Code LLaMA & Code LLaMA Instruct (Rozière et al., 2023). Code Llama is a code-specialized version of Llama2 (Touvron et al., 2023b) trained on code-specific datasets. Code LLaMA Instruct is a version further fine-tuned with human instructions and self-instruct code synthesis.
  • StarCoder (Li et al., 2023b). StarCoder is a widely-used large code language model trained on diverse sources, including 80+ programming languages, Git commits, GitHub issues, and Jupyter notebooks. It’s also one of the foundation models in our paradigm experiments.
  • WizardCoder (Luo et al., 2023). WizardCoder is instruction fine-tuned on StarCoder with 78k instruction data generated through the application of Code Evol-Instruct.
The selected models for comparison in NLU tasks include:
  • LLaMA (Touvron et al., 2023a). LLaMA is a collection of foundation language models trained on trillions of tokens from publicly available datasets.
  • LLaMA2 (Touvron et al., 2023b). Llama 2 is an updated version of Llama, trained on a new mix of publicly available data. It increased the size of the pretraining corpus by 40%, doubled the context length of the model, and adopted grouped-query attention in training.
  • Mistral Jiang et al. (2023). Mistral is a state-of-the-art 7B foundational model, fast-deployed, easily customizable, and supports English and code with an 8k context length. It’s also one of the foundation models in our paradigm experiments.
  • Vicuna Chiang et al. (2023). Vicuna is an open-source chatbot trained by fine-tuning LLaMA on 70K user-shared conversations collected from the ShareGPT website.
  • WizardLM Xu et al. (2023). WizardLM is instruction fine-tuned on LLaMA with 70k instruction data generated through the Evol-Instruct strategy.
  • Platypus (Lee et al., 2023). Platypus is a family of fine-tuned and merged LLMs achieving strong performance. It uses Open Platypus as its instruction dataset and applies the LoRA strategy to train adaptors that can be merged into different foundation models, creating many variant models.